every time the pedals go through a 360 degree rotation on a cetain bicycle, the tires rotate three times. if the tires are 24 inches in diameter, what is the minimum number of complete rotations of the pedals needed for the bicycle to travel at least one mile?
1) 12 3)561
2) 281 4)5,280

Answer by mananth(16949)

The diameter of the tyre is 24 inches
radius = 12 inches.
the circumference = 2*pi*r
=2*pi*12
=75.4 inches
If the cycle has to go 1 mile it has to go 63360 inches
the number of revolutions the tyre has to make = 63360 / 75.4
= 840.31 revolutions
if the tyre makes 840.31 revolutions the the pedal has to make
840.31 / 3
= 280.1 revolutions